### v1

# Scatterless interferences: Delay of laminar-to-turbulent flow transition by a lattice of subsurface phonons

Abstract
Turbulent flows in boundary layers generate significantly higher surface friction drag compared to laminar flows. Avoidance or delay of laminar-to-turbulent flow transition is therefore critical for efficient operation of air vehicles and water vessels. Here we present a passive phonons-based solution that is based on engineered intervention with the unstable flow fluctuations that trigger the transition. The intervention is introduced by a tailored frequency- and wavevector-dependent interaction between the flow and the bounding surface. The surface is homogenous and smooth with all the engineered constituents buried in the subsurface. The subsurface configuration comprises a wall-parallel lattice of wall-normal phononic structural units. Each discrete phononic subsurface unit is designed to respond to the flow in a negative phase as to inhibit the flow instabilities at the point of contact. The collective effect of the lattice formation of these units, in turn, offers a form of interferences that causes the kinetic energy of the flow instabilities to decay downstream to the location of the intervention, thus delaying the transition. This concept presents a new paradigm for the design of aircrafts and ships beyond the prevailing paradigm of streamlined shaping of the surfaces.
The interaction of a flowing fluid with a surface is a dynamical process that lies at the heart of a wide range of both natural and artificial systems. Two most consequential examples are the flow over an aircraft wing or around the hull of a ship. The nature of the fluid-structure interaction in these cases dictates the intensity of skin-friction drag, the location of onset of laminar-to-turbulent flow transition, and the possibility of flow separation--all with far-reaching implications on the overall performance of the vehicle [1, 2]. Skin friction drag in particular has profound economic and environmental costs because it directly affects the fuel efficiency. Effective intervention that delays flow transition is thus critical because wall-bounded turbulent flows generate significantly higher surface friction drag compared to laminar flows. Flow separation, as a contrasting scenario, is key to overall aerodynamic/hydrodynamic stability, form drag, vehicle maneuverability, and applications that require turbulent mixing. Similar considerations for fluid-structure interaction are applicable to other problems, such as land vehicles, wind turbines, and long-range pipelines, to name a few.
Laminar-to-turbulent transition in walled-bounded flows has been investigated extensively in the past, leading to a rich build-up of knowledge on how transition occurs across various flow regimes [3]. A “natural” form of transition is associated with the growth in amplitude of infinitesimal fluctuations (also referred to as perturbations or disturbances  [^1] These terms are used interchangeably in this article.) representing unstable propagating modes inherent in the flow. If left without intervention, these modes advance to nonlinear growth, secondary transition mechanisms, flow structure breakdown, and finally fully developed turbulent flow [3]. A dominant primary mode in air or water channel flows or boundary layers is known as Tollmien-Schlichting (T-S) waves [5, 6]. These unstable waves, which travel with the mean flow, have a vortical nature and grow over an identifiable narrow band of frequencies that can be predicted with linear stability analysis [7, 8]. Although T-S waves are not always the mechanism directly responsible for transition, particularly under realistic flight conditions, much research has been devoted to their understanding [9] due to their fundamental nature, and because they serve as a platform for development of new technologies for laminar flow control involving more complex types of instabilities.
Figure 1: Illustration of scatterless interferences: (a) Classical scattering of acoustic waves by a lattice of elastic objects, where the interfering waves and the objects are in the same space. (b) Scatterless interferences of flow fluctuation waves by a lattice of PSubs, where the waves and the PSubs are not in the same space.
Numerous research approaches have been proposed in the last couple of decades aimed at passively mitigating the undesirable effects of fluid-structure interactions. These include installation of riblets over the surface [10, 11], creation of artificial surface roughness [12, 13], drilling of holes to form a porous surface [14, 15], or coating the surface with a compliant (low stiffness) material [16, 17, 18, 19, 20, 21, 22]. None of these approaches, however, synchronize with the frequency, phase, and wavevector characteristics of the flow instabilities, and are therefore limited in their effectiveness. An ideal intervention requires a tailored solution, with mechanistic precision, to create a passive and responsive control stimulus that accounts for the dynamical properties of the underlying flow transition mechanisms. In 2015, the general concept of flow control by subsurface phonons has been introduced as an approach capable of achieving this level of precise wave-synchronized control of flow instabilities [23]. A phononic subsurface (PSub) comprises a synthetically designed architectured material affixed beneath  [^2] Installation of a PSub may be effectively concealed because it is placed underneath the surface. In other words, an external observer viewing a PSub-installed airplane wing would view a normally looking surface and would not be able to identify if a PSub is present or not. the surface exposed to the flow (e.g., of a wing or vehicle body). The function of the PSub is to manipulate small-amplitude vibrations on the surface, and by extension the flow fluctuations near the wall that are responsible for transition. PSubs may be designed to passively respond to the flow instabilities in an out-of-phase manner, creating stabilization, or in an in-phase manner, creating destabilization--either function is realized by design. In past computational investigations, a PSub has been applied as a solitary unit [23, 25, 26, 27, 28] or as a contiguous layout of units distributed along the streamwise direction [29, 30].
Despite the rapid progress of research on flow control by subsurface phonons, critical aspects remain to be addressed for the general concept to reach its potential. Two key limitations of previous demonstrations are that the PSub effect is not effective downstream to the PSub location, and that it is applicable only on unidirectional instability waves. The former must be resolved to enable transition delay, and the latter is significant for real-world flow control. While the downstream control objective was addressed with a “multiple-input-multiple-output” PSub configuration, offering a remarkable explicit display of transition delay [25], that approach is inherently limited to unidirectional instabilities because it dictates that a single PSub must interact with more than one flow point requiring a phased relation to be tuned specifically along a certain direction. Furthermore, the fixed distance between the input/output points limits the approach to a narrow band of fluctuation wavenumbers. Schmidt et al. [28] explored the concept from an alternative angle: rather than engage with subsurface phonons to inhibit the fluctuation production mechanisms (as done in all previous PSub studies), they designed their PSub based on pass-band motion to absorb and trap the fluctuation energy by resorting to actively controlled time-varying material properties. This approach enhances the energy transfer from the mean flow to the instabilities, rather than reduces it [26]. When more energy is drawn from the mean flow, this energy will accumulate in the subsurface structure requiring its dissipation as heat.
In this work, we present the concept of a lattice of PSubs comprising a collection of individual PSub units laid out following a square or hexagonal lattice symmetry  [^3] Other lattice symmetries may also be considered in the future.. Similar to acoustic or elastic waves propagating around or through a lattice of scatterers (see Fig. 1a), we design our configuration on the basis of a rigorous Bloch wave analysis where the field variable is the flow fluctuations. Yet, in our system there are no scatterers. In contrast to classical scattering problems where the interfering waves and the scattering objects are in the same space, the flow instability waves interfere as a result of the local influence of each PSub on the fluctuation velocity components. Here we recall that the PSub is beneath the space of the flow (see Fig. 1b). After developing Bloch’s theorem for this unique problem, we demonstrate by direct numerical simulations of the coupled fluid-Psub lattice system sustained downstream reduction in the fluctuations kinetic energy, which indicates a delay in laminar-to-turbulent transition. Moreover, as dictated by the PSub lattice symmetry, this approach is immune to any changes in the direction of propagation of the instability waves and may be tuned to accommodate a wide range of wavenumbers along each direction--similar to the classical acoustic Bragg scattering problem which we also demonstrate to provide a direct analogy and comparison.
Results
We form our lattice of Psubs using identical PSub units, all installed in a channel model. We select a channel flow problem for simplicity, but all the underlying concepts are readily applicable to boundary layer flows. The Psub design used in this investigation is shown in Fig. 2a. It consists of an elastic rod, composed of five unit cells, with each unit cell accomodating a local resonator (which in practice may be realized as a cantilevered pillar). The dispersion curves and the amplitude and phase frequency-response characteristics for this nominal PSub configuration is shown in Fig. 2b. The product of the amplitude and phase yields a performance metric which provides an a priori prediction of the beahvior of the PSub once passively engaged with the flow--this quantity is also plotted in Fig. 2b.
Figure 3 shows the results of applying Bloch analysis over a unit cell comprising the flow interacting with the PSub in three configurations, a full-span PSub, and the square and hexagonal lattice arrangements [see schematics (b), (c) and (d) in Fig. 3]. The PSub interaction with the flow has been accounted for through a frequency-dependent complex-valued admittance, denoted by $Y$ in Fig. 3(a). The admittance is obtained, for a given frequency, from the solution of the PSub system of equations [25] and it is closely related to the performance metric. We use this function alongside transpiration boundary conditions to find the most unstable eigenvalue--and associated eigenmode--from the standard Orr-Sommerfeld equations modified to accommodate the periodicity imposed by the application of Bloch’s theorem in terms of the T-S wave’s wavenumber, which corresponds to the eigenvalue. The mathematical details of the formulation can be found in the Supplementary Material. As a result of this analysis, the dispersion curves corresponding to the unstable flow mode in Fig. 3(a) have been obtained. These results are complemented by the rate of production of perturbation kinetic energy (PKE)--denoted by $P_{\text{r}}$ in Fig. 3--obtained from the Bloch analysis. These have been computed from the eigenmodes associated with the unstable eigenvalues, normalized by the average PKE computed for the rigid case (refer to Supplementary Materials for details). Results for both weak destabilizing and stabilizing cases, at frequencies of 1539.5 Hz and 1555 Hz respectively, have been averaged over a flow domain close to the PSub wall, in particular along 10% of the channel’s half height. Contour plots showing the $P_{\text{r}}$ spanwise distribution over three unit cells (in the streamwise direction) are given in Fig. 3(b) to (d) for the full-span and for each of the lattice configurations. The average $P_{\text{r}}$ values over a spanwise period are shown in the last row of Fig. 3(b) to (d) to summarize the results. The different cases are compared with the results for the rigid case, to show the spatial stabilizing/destabilizing effects when the flow interacts with each PSub patch.
The ultimate test for any flow control approach is through high-fidelity simulations, particularly direct numerical simulations (DNS) of the Navier-Stokes equations, as these resolve the high spatial and temporal frequencies of the flow fluctuations--which are central to the transition mechanism. We execute simulations in a three-dimensional (3D) computational domain and retain all nonlinear terms. In Fig. 4 we present DNS results for the two lattice configurations, and also show the results of an analogous acoustic scattering simulation for comparison. The coupled fluid-structure simulations were run at $Re=7500$ with a 1555 Hz unstable T-S wave input at the left end of the channel. See the Models and Methods sections for more details on the model and simulation parameters and numerical techniques used. The results in Figs. 4b and 4c show a sustained reduction in the PKE downstram of the PSub lattice, with a slightly stronger effect by the hexagonal arrangement compared to the square arrangement.
Discussion
The performance metric for the designed PSub unit shown in Fig. 2b shows that the unstable wave frequency of 1555 Hz falls at a negative value, indicative of a stabilization effect. While the surrounding negative region is relatively narrow--since it is associated with a given structural resonance of the finite PSub structure, future work will explore strategies to broaden the frequency range of operation.
Results from the Bloch analysis show a correlation between the dispersion curves (in Fig. 3) and the PSub performance metric (in Fig. 2). It is observed that the intensity of the T-S wave’s instability--in this case given by the imaginary wavenumber--rapidly grows in the vicinity of the resonance frequency, which helps enhance the stabilizing/destabilizing effects. Furthermore, the rate of production of PKE within the PSub region consistently increases in the destabilizing case and decreases--even reaching negative values--in the stabilizing case, in agreement with the stand-alone performance metric predictions (see Fig. 3(b) to (d)). Interestingly, the $P_{\text{r}}$ also exhibits an opposite effect upon leaving the PSub, which is in accordance with what we observe in the fluid-structure interaction (FSI) simulations (see, for instance, how the PKE slightly increases in Figs. 4(b) and (c) each time the flow leaves a PSub unit behind despite being stabilizing cases). This analysis also shows how the full-span case exhibits a stronger response than both lattice configurations, which demonstrates that spanwise interactions can affect the flow instabilities downstream, as demonstrated by the full-scale FSI simulations results. While both square and hexagonal PSub lattices are effective in sustaining downstream stabilization, both the Bloch analysis and simulations indicaete that the hexagonal configuration is slightly superior--its PKE slope in Fig. 4c is $\theta=4.6^{\circ}$ compared to $\theta=4.4^{\circ}$ for the square lattice. Upon further design optimization of both the PSub unit and the lattice configuration, stronger downstream reductions in PKE are attainable. It is clearly observed that both flow fluctuations “scatterless” interferences are nearly identical to the interferences triggered by acoustic scattering.
Figure 2: PSub design and its dispersion and vibration response characteristics: (a) Schematic of locally-resonant elastic metamaterial-based PSub utilized in this study. The length of the unit cell is $a_{\rm Psub}=1$ cm yielding a total PSub length of 5 cm. Each PSub is installed in the flow subsurface with its top face directly exposed to the flow. Flow instabilities, e.g. T-S waves, will continuously excite the PSub at the top edge and the PSub, in turn, will respond at the same point. Depending on the frequency, the PSub will respond according to its dynamical characteristics. Close to an anti-resonance or a resonance, the PSub motion will respectively impede or enhance the energy extraction from the mean flow into the fluctuation field [26]. This passive process will repeat and cause sustained control of incoming instability waves near the wall at the location of the PSub. (b) Four key plots that characterize the PSubs dynamics. From top to bottom, the following are shown: dispersion curves for PSub unit cell, steady-state vibration amplitude and phase response of the PSub top edge when excited at the same location, and performance metric obtained by multiplying the amplitude by the phase. All plots are obtained by analyzing a stand-alone finite-element model of the PSub.
Figure 3: (a) Dispersion curves for selected unstable flow mode in the presence of a PSub lattice. The first and second rows correspond to the imaginary and real components of the wavenumber $\kappa$ for the rigid cases and the PSub cases in the full-span, square-lattice and hexagonal-lattice configurations [depicted by the (b), (c) and (d) unit cell schematics, respectively]. The third row shows the real and imaginary components of the PSub’s admittance $Y$ at different frequencies. The dashed vertical lines indicate the frequencies for which the average modal perturbation kinetic energy (PKE) production rate, $P_{\text{r}}$, distributions on the right plots have been obtained. Results for each frequency are depicted in their corresponding panel. The first panel (in orange) shows a weak destabilizing case (at a frequency of 1539.5 Hz) and the second panel (in light green) is for a weak stabilizing case (at a frequency of 1555 Hz). The curves below the panels correspond to the averaged production rate evolution obtained for both the weak destabilizing and stabilizing cases compared to that of the rigid case. As a reference, the average has been evaluated in the vicinity of the PSub wall (in the range of 10% of half of the channel’s height), and all the modal amplitudes have been normalized to make them coincide with that of the rigid case at the left edge of the first unit cell.
Figure 4: Demonstration of scatterless interferences by DNS: (a) Acoustic wave propagation in a space covered by a lattice of elastic scatterers, considering square symmetry (left) and hexagonal symmetry (right). (b) Time-averaged flow PKE as a function of the streamwise and spanwise directions for flow in a channel with PSubs laid out following both symmetries. (c) Time- and spanwise-averaged PKE as a function of the streamwise position. The results in (b) and (c) indicate interference patterns, even though the surface is flat and smooth and there are no objects in the flow space. These interferences stem for the collective action of the local effect by each PSub. Figure 4(c) shows time-averaged PKE on and off the centerline of the channel, reflecting the effects of the different symmetries of the square and hexagonal lattices. The time- and spanwise-averaged PKE is shown in dark green, where the angle $\theta$ at the end of the channel indicates a decrease in the slope of PKE compared to the rigid case, and hence transition delay downstream of the PSub lattice. The angle is measured as $\theta=4.4^{\circ}$ and $\theta=4.6^{\circ}$ for the square and hexagonal cases, respectively, indicating the transition is delayed further by the hexagonal PSubs lattice.
In conclusion, we have demonstrated that by engineering the phonon properties of the subsurface, which could very well be concealed, it is possible to almost completely execute full command of the very nature of the fluid-structure interaction. Such intervention, which is fully passive, offers precise mechanistic control of the behavior of wall-bounded flows. Our findings demonstrate that the principles of Bloch analysis, traditionally applied to acoustic and elastic waves, can be effectively translated into the realm of flow control through the implementation of PSubs. By leveraging locally resonant metamaterial-based phononic subsurfaces, we have shown that it is possible to manipulate flow instabilities in a manner analogous to wave propagation in periodic acoustic and elastic media. This novel approach enables passive control of phenomena such as transition to turbulence, with lattice arrangements of PSubs facilitating attenuation of T-S waves’ instabilities downstream, while also being robust to variations in wavenumber and direction of propagation. With the advent of concepts from phonon engineering and metamaterials, the prospects of future improvements of the performance is highly accessible. It is conceivable, for example, that with architectured materials, a fully laminar boundary layer around an airplane may be attained in the near future. Such outcome would drastically improve fuel efficiency. Potentially, the Psub lattices may be designed to simultaneously inhibit noise emission as well, which would bring about the added advantage of passenger comfort and reduced noise pollution.
Models and Methods
[…]
Each PSub unit is modeled as a finite linear elastic metamaterial consisting of 5 rod unit cells with one local mass-spring resonator in the center of the unit cell [26, 30]. The PSub is free to deform at the edge interfacing with the flow (top) and is fixed at the other end (bottom). Each individual PSub is allowed to deform independently from the adjacent rigid wall and from the motion of neighboring PSubs; its top surface deformation takes a uniform profile across the fluid-PSub interface region [30]. The length of the unit cell along the wall-normal direction is $L_{\mathrm{UC}}=1$ $\mathrm{cm}$ (i.e., total PSub length is $5$ $\mathrm{cm}$). The resonator frequency are set to $\Omega_{\mathrm{res}}=2000$ $\mathrm{Hz}$ by tuning the resonators point masses to be ten times heavier than the total mass of the unit-cell base ($m_{\mathrm{res}}=10\times\rho L_{\mathrm{UC}}$), where $\rho$ is the base material density. Hence, the stiffness of the resonator spring is $k_{\mathrm{res}}=m_{\mathrm{res}}(2\pi f_{\mathrm{res}})^{2}$. The base is composed of ABS polymer with density of $\rho=1200$ $\mathrm{kg/m^{3}}$ and Young’s modulus of $E=3$ $\mathrm{GPa}$. Material damping is modeled as viscous proportional damping with constants $q_{1}=0$ and $q_{2}=6\times 10^{-8}$ [30].
The Navier-Stokes equations are integrated using a time-splitting scheme [34, 35, 36] on a staggered structured grid system. A two-node iso-parametric finite-element model is used for determining the PSub nodal axial displacements, velocities, and accelerations [40] where time integration is implemented simultaneously with the flow simulation using an implicit Newmark algorithm [41]. Since the equations for the fluid and the PSub are inverted separately in the coupled simulations, a conventional serial staggered scheme [42] is implemented to couple the two sets of time integration. More details on the computational models and numerical schemes used are detailed in Refs. [26, 30]. The relative geometric dimensions of the PSubs forming each of the square and hexagonal lattices are to scale as shown in Fig. 4, with more information available in the Supplemntal Materials document.
Acknowledgments
This work utilized the Summit supercomputer, which is supported by the National Science Foundation (awards ACI-1532235 and ACI-1532236), the University of Colorado Boulder, and Colorado State University. The Summit supercomputer is a joint effort of the University of Colorado Boulder and Colorado State University. This work also utilized the Alpine high performance computing resource at the University of Colorado Boulder. Alpine is jointly funded by the University of Colorado Boulder, the University of Colorado Anschutz, and Colorado State University and with support from NSF grants OAC-2201538 and OAC-2322260.