### v3

# Scatterless interference: Delay of laminar-to-turbulent flow transition by a lattice of subsurface phonons

Abstract
Wave interference has historically relied on scattering objects placed within the wave domain. Here, we introduce a fundamentally new mechanism: scatterless interference induced by a lattice of subsurface phonon motion beneath a smooth wall interfacing with a transitioning boundary-layer flow. The subsurface consists of a wall-parallel lattice of wall-normal frequency-dependent phononic structural units, each designed to respond to local flow perturbations in an out-of-phase manner, suppressing them at the point of interaction. Collectively, the lattice induces interference effects that cause the kinetic energy of flow instabilities to decay downstream, thereby delaying laminar-to-turbulent transition. To guide the design of the phononic subsurface lattice, a Bloch-wave unit-cell analysis is developed for the flow perturbations, and direct numerical simulations validate the concept. This work establishes scatterless interference as a distinct physical phenomenon and represents a paradigm shift in the design of aerodynamic and hydrodynamic surfaces—moving beyond streamlined shaping to leveraging subsurface phonon engineering for drag reduction and enhanced performance.
Introduction
The control of wave interference through the deliberate placement of scattering objects has been a cornerstone principle in wave physics for over two centuries. Its classical origins trace back to Thomas Young’s double-slit experiment [1], which first demonstrated the wave nature of light via interference between scattered paths. This concept evolved into a general principle applied across multiple disciplines. In electromagnetics, the development of photonic crystals [2, 3] relied on arrays of dielectric scatterers directly interacting with the wave field to trigger spatial constructive and destructive interferences that bring rise to Bragg scattering. A similar mechanism emerged in acoustics and elasticity, where periodic voids or inclusions, forming phononic crystals [4, 5], have enabled the formation of wave interference patterns and spectral band gaps through in-domain scattering. Even in quantum mechanics, interference effects typically arise through the interaction of wavefunctions with potentials or obstacles embedded in the same domain of motion. These classical and modern mechanisms share a foundational attribute: the structures responsible for interference are located within the same spatial domain as the wave field they manipulate  [^1] The only exception is having active external excitations to trigger the scattering [55, 56, 57]. To date, this paradigm has remained essentially unchallenged across physics. In this work, we introduce a fundamentally new mechanism of passive wave interference that breaks from the conventional paradigm: scatterless interference induced by a lattice of phononic subsurface units. In sharp contrast to prevalent scattering mechanisms, the interference patterns in our fully passive system arise not from scatterers embedded within the wave domain, but from elastic phononic structures buried beneath the surface—outside the spatial domain occupied by the propagating waves. We present this new physical mechanism in the context of the long-standing classical fluid dynamics problem of laminar-to-turbulent transition.
Fluid-structure interaction
The laminar-to-turbulent transition in wall-bounded flows has been the subject of extensive research, resulting in a rich build-up of knowledge on how transition unfolds across different flow regimes [7]. A “natural” transition pathway typically involves the growth of infinitesimal flow perturbations--also referred to as disturbances  [^2] These terms are used interchangeably in the literature.--which represent inherent unstable modes within the flow. Without intervention, these modes may advance to nonlinear amplification, leading to flow structure breakdown and the ultimate evolution into fully developed turbulence [7]. A dominant primary mode in air or water channel flows, as well as boundary layers, is known as the Tollmien-Schlichting (T-S) wave [9, 10]. Unstable T-S waves take a vorticial form and travel with the mean flow, growing over an identifiable narrow band of frequencies that can be predicted with linear stability analysis [11, 12]. While T-S waves are not always the direct cause of transition, especially under complex or realistic operating conditions, they have been the focus of extensive research [13]. This is due to their fundamental nature, and because they serve as a platform for development of new technologies for laminar flow control involving more complex types of instabilities.
Figure 1: Illustration of scatterless interferences: (a) Classical scattering of acoustic waves by a lattice of elastic objects, where the interfering waves and the objects are in the same space. (b) Scatterless interferences of flow perturbations waves by a lattice of PSubs, where the waves and the PSubs are not in the same space.
Numerous research approaches have been proposed in the last couple of decades aimed at passively mitigating the undesirable effects of fluid-structure interactions. These include installation of riblets over the surface [14, 15], creation of artificial surface roughness [16, 17], drilling of holes to form a porous surface [18, 19], or coating the surface with a compliant (low-stiffness) material [20, 21, 22, 23, 24, 25, 26]. These approaches, however, are not based on a synchronization with the frequency, phase, and wavevector characteristics of the flow instabilities, and are therefore limited in their effectiveness. An ideal intervention requires a tailored solution, with mechanistic precision, to create a passive and responsive control stimulus that accounts for the dynamical properties of the underlying flow transition mechanisms. In 2015, the general concept of flow control by subsurface phonons was introduced as an approach capable of achieving this level of precise wave-synchronized control of flow instabilities [27]. A phononic subsurface (PSub) comprises a synthetically designed architected material affixed beneath  [^3] Installation of a PSub may effectively be concealed because it is placed underneath the surface. In other words, an external observer viewing a PSub-installed airplane wing would view a normally looking surface and would not be able to identify if a PSub is present or not. the surface exposed to the flow (e.g., of a wing or vehicle body). The function of the PSub is to manipulate small-amplitude vibrations on the surface, and by extension the flow perturbations near the wall that are responsible for transition. PSubs may be designed to passively respond to the flow instabilities in an out-of-phase manner, creating stabilization, or in an in-phase manner, creating destabilization--either function is realized a priori by design, or in the future may be switched or tuned in real time by actuation. In past computational investigations, a PSub has been applied as a solitary unit [27, 29, 30, 31, 32] or as a contiguous layout of units distributed along the streamwise direction [33, 34].
Despite the rapid progress of research on flow control by subsurface phonons, critical aspects remain to be addressed for the general concept to reach its potential. Two key limitations of previous demonstrations are that the PSub effect is not effective downstream to the PSub location, and that it is applicable only to unidirectional instability waves. The former must be resolved to enable transition delay, and the latter is significant for real-world flow control--where cross-flow scenarios may be encountered, for example. While the downstream control objective was addressed with a “multiple-input-multiple-output” PSub configuration, offering a remarkable explicit display of transition delay [29], that approach is inherently limited to unidirectional instabilities because it dictates that a single PSub must interact with more than one flow point requiring a phased relation to be tuned specifically along a certain direction. Furthermore, the fixed distance between the input/output points limits the approach to a narrow band of perturbation wavenumbers. Schmidt et al. [32] explored the concept from an alternative angle: rather than engage with subsurface phonons to inhibit the perturbation production mechanisms (as done in all previous PSub studies), they designed their PSub based on pass-band motion to absorb and trap, by filtering, the energy of undesirable fluctuations. Actively controlled time-varying material properties were used for demonstration, however the principle may be implemented passively using nonlinear effects. This approach, which is applicable to only the function of stabilization, permanently attracts the energy that is continuously being transferred from the mean flow to the instabilities, instead of reducing it at the source [30]. This energy will accumulate in the subsurface structure requiring its ultimate dissipation as heat.
Scatterless interference
In this work, we present the concept of a lattice of PSubs comprising a collection of individual PSub units laid out following a square or hexagonal lattice symmetry  [^4] Other lattice symmetries may also be considered in the future.. Similar to acoustic or elastic waves propagating around or through a lattice of rigid or elastic scatterers, respectively (see Fig. 1a), we design our configuration on the basis of a rigorous Bloch wave analysis where the field variable is the flow perturbations. Yet, in our system there are no scatterers. In contrast to classical scattering problems where the interfering waves and the scattering objects are located in the same space, the flow instability waves interfere as a result of the local influence of each PSub on the perturbation velocity components. Here we recall that the PSubs are located beneath the spatial domain of the flow (see Fig. 1b). After developing Bloch’s theorem for this unique problem, we demonstrate by direct numerical simulation (DNS) of the coupled fluid-PSub lattice systems sustained downstream reduction in the perturbation kinetic energy (KE), which indicates a delay in laminar-to-turbulent transition. Moreover, as dictated by the PSub lattice symmetry, this approach is immune to any changes in the direction of propagation of the instability waves and may be tuned to accommodate a relatively wide range of frequencies and wavenumbers along each direction--similar to the classical acoustic Bragg scattering problem which we also demonstrate to provide a direct analogy and comparison.
Figure 2: PSub design and its dispersion and vibration response characteristics: (a) Schematic of locally-resonant elastic metamaterial-based PSub unit utilized in this study. The length of the unit cell is $a_{\rm Psub}=1$ cm yielding a total PSub length of 5 cm. Each PSub is installed in the flow subsurface with its top face directly exposed to the flow. Flow instabilities, e.g. T-S waves, will continuously excite the PSub at the top edge and the PSub, in turn, will respond at the same point. Depending on the frequency, the PSub will respond according to its dynamical characteristics. Following an out-of-phase or in-phase response, the PSub motion will respectively impede or enhance the energy extraction from the mean flow into the perturbation field [30]. This passive process will repeat and cause sustained control of incoming instability waves near the wall at the location of the PSub. (b) Four key plots that characterize the PSubs dynamics. From top to bottom, the following are shown: dispersion curves for PSub unit cell, steady-state vibration amplitude and phase response of the PSub top edge when excited at the same location, and performance metric obtained by multiplying the amplitude by the phase. All plots are obtained by analyzing a stand-alone finite-element model of the PSub.
Results
We form our lattice of PSubs using identical PSub units, all installed in a channel model. We select a channel flow problem for simplicity, but all the underlying concepts are readily applicable to boundary layer flows. The design of the PSub unit used in this investigation is shown in Fig. 2a. It consists of an elastic rod, composed of five unit cells, with each unit cell accommodating a local resonator (which in practice may be realized as a cantilevered beam or pillar). The dispersion curves and the amplitude and phase frequency-response characteristics for this nominal PSub configuration is shown in Fig. 2b. The product of the amplitude and phase yields a performance metric [27] which provides an a priori prediction of the behavior of the PSub once passively engaged with the flow--this quantity is also plotted in Fig. 2b.
Figure 3: (a) Dispersion curves for selected unstable flow mode in the presence of a PSub lattice. The first and second rows correspond to the imaginary and real components of the wavenumber $\kappa$ for the rigid-wall case and the PSub cases in the full-span, square-lattice, and hexagonal-lattice configurations [depicted by the (b), (c) and (d) unit cell schematics, respectively]. The third row shows the real and imaginary components of the PSub’s admittance $Y$ at different frequencies. The dashed vertical lines indicate the frequencies for which the average modal perturbation KE production rate, $P_{\text{r}}$, distributions on the right plots have been obtained. Results for each frequency are depicted in their corresponding panel. The first panel (in orange) shows a weak destabilizing case (at a frequency of 1539.5 Hz) and the second panel (in light green) is for a weak stabilizing case (at a frequency of 1555 Hz). The curves below the panels correspond to the averaged production rate evolution obtained for both the weak destabilizing and weak stabilizing cases compared to that of the rigid case. As a reference, the average has been evaluated in the vicinity of the PSub wall (in the range of 10% of half of the channel’s height), and all the modal amplitudes have been normalized to make them coincide with that of the rigid-wall case at the left edge of the first unit cell.
Figure 4: Demonstration of scatterless interferences of flow perturbations by DNS: (a) Acoustic wave propagation in a space covered by a lattice of elastic scatterers, considering square symmetry (left) and hexagonal symmetry (right). (b) Time-averaged flow perturbation KE as a function of the streamwise and spanwise directions for flow in a channel with PSubs laid out following both symmetries. (c) Time- and spanwise-averaged perturbation KE as a function of the streamwise position. The results in (b) and (c) indicate interference patterns, even though the surface is flat and smooth and there are no objects in the flow space. These interferences stem for the collective action of the local effect by each PSub. Figure 4(c) shows time-averaged perturbation KE on and off the centerline of the channel, reflecting the effects of the different symmetries of the square and hexagonal lattices. The time- and spanwise-averaged perturbation KE is shown in dark green, where the angle $\theta$ at the end of the channel indicates a decrease in the slope of $K_{\text{p}}$ compared to the rigid-wall case, and hence transition delay downstream of the PSub lattice. The angle is measured as $\theta=4.4^{\circ}$ and $\theta=4.6^{\circ}$ for the square and hexagonal cases, respectively, indicating the transition is delayed further by the hexagonal PSubs lattice.
Figure 3 shows the results of applying Bloch analysis over a unit cell comprising the flow interacting with the PSub in three configurations, a full-span PSub, and the square and hexagonal lattice arrangements [see schematics (b), (c) and (d) in Fig. 3]. The PSub interaction with the flow has been accounted for through a frequency-dependent complex-valued admittance, denoted by $Y$ in Fig. 3(a). The admittance is obtained, for a given frequency, from the solution of the PSub system of equations [29] and it is closely related to the performance metric. We use this function alongside transpiration boundary conditions to find the most unstable eigenvalue--and associated eigenmode--from the standard Orr-Sommerfeld equations modified to accommodate the periodicity imposed by the application of Bloch’s theorem in terms of the T-S wave’s wavenumber, which corresponds to the eigenvalue. The mathematical details of the formulation and definitions of key quantities can be found in the Supplementary Material. As a result of this analysis, the dispersion curves corresponding to the unstable flow mode in Fig. 3(a) have been obtained. These results are complemented by computing the average perturbation kinetic energy, denoted by $K_{\text{p}}$, and the rate of production of that energy, denoted by $P_{\text{r}}$. The perturbation KE has been computed from the eigenmodes associated with the unstable eigenvalues, normalized by the corresponding value for the rigid case. The $P_{\text{r}}$ quantity, as obtained from the Bloch analysis, is plotted in Fig. 3b-c. The reader is referred to Supplementary Materials for calculation details. Results for both a weak destabilizing case and a weak stabilizing case, at frequencies of 1539.5 Hz and 1555 Hz respectively, have been averaged over a flow domain close to the PSub wall, in particular along 10% of the channel’s half height. Contour plots showing the $P_{\text{r}}$ spanwise distribution over three unit cells (in the streamwise direction) are given in Fig. 3(b) to (d) for the full-span and for each of the lattice configurations. The average $P_{\text{r}}$ values over a spanwise period are shown in the last row of Fig. 3(b) to (d) to summarize the results. The different cases are compared with the results for the rigid case, to show the spatial stabilizing/destabilizing effects when the flow interacts with each PSub unit.
The ultimate test for any flow control approach is through high-fidelity simulations, particularly DNS of the Navier-Stokes equations, as these resolve the high spatial and temporal frequencies of the flow perturbations--which are central to the transition mechanism. We execute simulations in a three-dimensional (3D) computational domain and retain all nonlinear terms. In Fig. 4 we present DNS results for the two lattice configurations, and also show the results of an analogous acoustic scattering simulation for comparison. The coupled fluid-structure simulations were run at $Re=7500$ with a 1555 Hz unstable T-S wave input at the left end of the channel. See the Models and Methods sections for more details on the model, simulation parameters, and numerical techniques used. The results in Figs. 4b and 4c show a sustained reduction in $K_{\text{p}}$ downstream of the PSub lattice, with a slightly stronger effect by the hexagonal arrangement compared to the square arrangement. This behavior is indicative of a delay in laminar-to-turbulent transition downstream of the PSub lattice region.
Discussion
The performance metric for the designed PSub unit shown in Fig. 2b shows that the unstable wave frequency of 1555 Hz falls at a negative value, indicative of a stabilization effect. While the surrounding negative region is relatively narrow--since it is associated with a given structural resonance of the finite PSub structure, future work will explore strategies to broaden the frequency range of operation.
Results from the Bloch analysis show a correlation between the dispersion curves (in Fig. 3) and the PSub performance metric (in Fig. 2). It is observed that the intensity of the T-S wave’s instability--in this case given by the imaginary wavenumber--rapidly grows in the vicinity of the resonance frequency, which helps enhance the stabilizing/destabilizing effects. Furthermore, the rate of production $P_{\text{r}}$ within the PSub region consistently increases in the destabilizing case and decreases--even reaching negative values--in the stabilizing case, in agreement with the stand-alone performance metric predictions (see Fig. 3(b) to (d)). Interestingly, $P_{\text{r}}$ also exhibits an opposite effect at the trailing edge of each PSub unit, which is in accordance with what we observe in the fluid-structure interaction (FSI) simulations (see, for instance, how the $K_{\text{p}}$ slightly increases in Fig. 4(b), for both lattice configurations, each time the flow “leaves” a PSub unit behind despite being stabilizing cases). This analysis also shows that the leading and trailing edge changes in $P_{\text{r}}$ are stronger in the full-span case compared to the two lattice configurations, except that this case does not experience spanwise interactions. While both square and hexagonal PSub lattices are effective in sustaining downstream stabilization, both the Bloch analysis and simulations indicate that the hexagonal configuration is slightly superior--its $K_{\text{p}}$ slope in Fig. 4c is $\theta=4.6^{\circ}$ compared to $\theta=4.4^{\circ}$ for the square lattice. Upon further design optimization of both the PSub unit and the lattice configuration, stronger downstream reductions in $K_{\text{p}}$ are attainable. It is intriguingly observed that the flow perturbations “scatterless” interferences resemble the interferences triggered by acoustic scattering, for both lattice configurations.
In conclusion, the mechanism of scatterless interference reported here reveals a unique class of wave behavior that to our knowledge has not been previously reported in any branch of physics. The concept comprises subsurface phononic units arranged with full spatial flexibility and designed to vibrate at specific frequencies, enabling precise, tunable, and frequency-selective interference patterns in the overlying wave field. The physical platform we employ to demonstrate this mechanism is an unstable laminar flow field subject to perturbation waves—a canonical setting for studying aerodynamic or hydrodynamic instability. We have shown that by engineering the phonon properties of the subsurface--virtually entirely concealed beneath the flow-exposed surface--it is possible to exert full command over the very nature of the underlying fluid-structure interaction. This passive intervention offers precise mechanistic and tunable control of the behavior of wall-bounded flows in general. Furthermore, our findings demonstrate that the principles of Bloch wave analysis, traditionally applied to electronic, photonic, acoustic, and elastic waves [36, 37], can be effectively extended to flow perturbations influenced by PSubs. By leveraging locally resonant metamaterial-based individual PSub units, we have demonstrated that flow instabilities can be manipulated through interference effects, in a manner analogous to wave phenomena in periodic elastic or acoustic media. This approach enables passive control over critical processes such as the transition to turbulence, with the PSub lattice facilitating attenuation of unstable T-S waves downstream, while being robust to variations in wavenumber and direction of propagation. Other types of flow instabilities may be similarly treated in future studies. With the ongoing influx of new concepts from phonon engineering and metamaterials [38, 39, 40], the prospects of further future improvements of the performance following this framework is highly accessible, and so is the path to practical implementation given the rapid advancements in fabrication technology [41, 42]. A wide range of flow-control applications stand to benefit, spanning from hypersonic airflows to low-speed liquid transport. It is conceivable that within this framework future aircrafts will maintain a fully laminar boundary layer over their surfaces, enabling transformative improvements in fuel efficiency. In addition, PSub lattices may be simultaneously designed to inhibit aeroacoustic emissions, offering the additional rewards of improved passenger comfort and reduced environmental noise pollution. Similar advances will materialize for hydroacoustic emissions as well.
Models and Methods
[…]
Each PSub unit is modeled as a finite linear elastic metamaterial consisting of 5 rod unit cells with one local mass-spring resonator in the center of the unit cell [30, 34]. The PSub is free to deform at the edge interfacing with the flow (top) and is fixed at the other end (bottom). Each individual PSub is allowed to deform independently from the adjacent rigid wall and from the motion of neighboring PSubs; its top surface deformation takes a uniform profile across the fluid-PSub interface region [34]. The length of the unit cell along the wall-normal direction is $L_{\mathrm{UC}}=1$ $\mathrm{cm}$ (i.e., total PSub length is $5$ $\mathrm{cm}$). The resonator frequency is set to $\Omega_{\mathrm{res}}=2000$ $\mathrm{Hz}$ by tuning the resonator’s point mass to be ten times heavier than the total mass of the unit-cell base ($m_{\mathrm{res}}=10\times\rho L_{\mathrm{UC}}$), where $\rho$ is the base material density. Hence, the stiffness of the resonator spring is $k_{\mathrm{res}}=m_{\mathrm{res}}(2\pi f_{\mathrm{res}})^{2}$. The base is composed of ABS polymer with density of $\rho=1200$ $\mathrm{kg/m^{3}}$ and Young’s modulus of $E=3$ $\mathrm{GPa}$. Material damping is modeled as viscous proportional damping with constants $q_{1}=0$ and $q_{2}=6\times 10^{-8}$ [34].
The Navier-Stokes equations are integrated using a time-splitting scheme [45, 46, 47] on a staggered structured grid system. A two-node iso-parametric finite-element model is used for determining the PSub nodal axial displacements, velocities, and accelerations [51] where time integration is implemented simultaneously with the flow simulation using an implicit Newmark algorithm [52]. Since the equations for the fluid and the PSub are inverted separately in the coupled simulations, a conventional serial staggered scheme [53] is implemented to couple the two sets of time integration. This scheme has been extensively verified, yielding excellent agreement with the experimentally validated linear theory with a maximum deviation of 0.05% in the predicted perturbation energy growth [54].More details on the computational models and numerical schemes used are detailed in Refs. [30, 34]. The relative geometric dimensions of the PSubs forming each of the square and hexagonal lattices are to scale as shown in Fig. 4, with more information available in the Supplemntal Materials document.
Acknowledgments
This research project is currently being supported by Office of Naval Research Multidisciplinary University Research Initiative (MURI) Grant Number N0001421268. The presented work utilized the Summit supercomputer, which is supported by the National Science Foundation (awards ACI-1532235 and ACI-1532236), the University of Colorado Boulder, and Colorado State University. The Summit supercomputer is a joint effort of the University of Colorado Boulder and Colorado State University. This work also utilized the Alpine high performance computing resource at the University of Colorado Boulder. Alpine is jointly funded by the University of Colorado Boulder, the University of Colorado Anschutz, and Colorado State University and with support from NSF grants OAC-2201538 and OAC-2322260.